Complex Time Evolution of Simple Quantum Systems

Joel Lebowitz Rutgers

I will describe some exact results about the long time behavior of a simple quantum system, having both bound and continuum states, when subjected to a periodic external force. This turns out to be surprisingly complex.

Joint work with O. Costin, R. Costin and A. Rokhlenko.